.                                                                                                   

In 2019, 200 KGs of oranges are produced at Rs.40/kg, 500 KGs of apples are produced at 120/kg, and 200 Glasses of orange juice are produced at Rs.150/glass.

In 2020, 220 KGs of oranges are produced at Rs.50/kg, 450 KGs of apples are produced at 140/kg, and 230 glasses of orange juice are produced at Rs.135/glass.

Assuming 2019 is the base year; calculate the nominal GDP in 2019



1
Expert's answer
2021-05-07T09:55:16-0400

Nominal GDP = Sum of (this year prices x this year quantities)

Nominal GDP 2019:

Oranges "= 40 \times × 200 = 8,000"

Apples "= 120 \times × 500 = 60,000"

Orange juice "= 200 \times × 150 = 30,000"

Total "= 8,000 + 60,000 + 30,000 = 98,000"

"Nominal\space GDP\space 2019 = 98,000"